OpenStudy (anonymous):

You are decorating a cylindrical hat box to give to your grandmother as a gift. You want to paste family photos onto the curved part of the box. The height of the box is 5 inches and the radius is 6 inches. If none of the photos overlap, what is the area of photographs you need? Use3.14 for pi.

OpenStudy (anonymous):

i got 94.2 i was wondering if this is right?

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

How did you get that answer?

zepdrix (zepdrix):

Hmm I'm confused where your numbers are coming from cydni D: like that 8. \[\large A=(Circumference)(Height)\]\[\large A=(2 \pi r)(h)\]\[\large A=(2\pi\cdot6)(5)\]\[\large A=60\cdot \pi\]\[\large A=60(3.14)=?\]